Let X N(0, 9) have mean 0 and variance 9. Find the expected value of X2(X +1).
Answer #1

X~normal distribution

E(X)=0 ,V(X)=9

V(X)=E(X^2)-[E(X)]^2

9=E(X^2)-0

E( X 2) = 9

E(X^3)=0

# for normal distribution

mu_{2r+1}=0

Where r=0,1,2......

E[X^2(X+1)]=E[X^3+X^2]

  =E[X^3]+E[X^2]

=0+9

=9